# Exact WKB in all sectors I: Potentials with degenerate saddles

### v1

@@ Abstract
We explore the exact-WKB (EWKB) method through the analysis of Airy and Weber types, with an emphasis on the exact quantization of locally harmonic potentials in multiple sectors. The core innovation of our work lies in introducing a novel complexification approach to the energy parameter $u$, distinct from the common complexification of the (semi-classical) expansion parameter used in Borel summability. This new technique allows for continuous analytical continuation across different sectors of a potential while maintaining the exact quantization condition, even before median summation. By redefining the $A$-cycle above the potential barrier top, we ensure the quantization condition remains real and, by use of the Stokes automorphism and the median resummation, show that the resurgence structure is preserved across transitions between sectors. Furthermore, we extend the Weber-type exact-WKB method, offering exact estimates for quantum actions around all types of saddle points, generalizing previous results. Through the analysis of these quantum actions, we reveal the presence of $S$-duality, facilitating the exchange between perturbative and non-perturbative behaviors, and we conjecture the mapping of the P-NP relations between dual theories. Our study encompasses periodic and symmetric double-well potentials, demonstrating that the exact-WKB method captures intricate structures in quantum systems in all sectors, including multi-instanton contributions and the resurgence of quantum actions.
@@ 1 Introduction
The exact-WKB (Wentzel-Kramers-Brillouin) method BPV ; Voros1983 ; Silverstone ; AKT1 ; AKT2 ; Takei1 ; DDP2 ; DP1 ; Takei2 ; Kawai1 ; Takei3 ; AKT3 ; AKT4 ; Schafke1 ; Iwaki1 stands as one of the most powerful tools in the study of quantum mechanical systems. With being combined with Borel resummation, the exact-WKB (EWKB) method offers significant improvements over semi-classical approximations, particularly in quantum systems where both perturbative and non-perturbative effects play a role. In recent decades, the EWKB method has seen significant advancements, fueled by contributions from both physicists and mathematicians. These developments have led to a deeper understanding of the connection between the semi-classical analysis and the exact quantization conditions Sueishi:2020rug ; Sueishi:2021xti ; Kamata:2021jrs , revealing intricate structures such as Stokes phenomena and resurgence theory in quantum mechanical systems Alvarez1 ; Alvarez2 ; Alvarez3 ; Jentschura:2004jg ; Zinn-Justin:2004vcw ; Zinn-Justin:2004qzw ; Jentschura:2010zza ; Jentschura:2011zza ; Basar:2013eka ; Dunne:2013ada ; Dunne:2014bca ; Basar:2015xna ; Misumi:2015dua ; Behtash:2015loa ; Gahramanov:2015yxk ; Kozcaz:2016wvy ; Dunne:2016jsr ; Dunne:2016qix ; Fujimori:2016ljw ; Basar:2017hpr ; Fujimori:2017oab ; Sueishi:2019xcj ; Cavusoglu:2023bai ; Cavusoglu:2024usn ; Ture:2024nbi . The method has been instrumental in exploring non-perturbative effects involving multi-instanton contributions which appear in analysis of various quantum mechanical and field theoretical systems Zinn-Justin:1981qzi ; Zinn-Justin:1983yiy ; Unsal:2007vu ; Unsal:2007jx ; Shifman:2008ja ; Poppitz:2009uq ; Anber:2011de ; Poppitz:2012sw ; Misumi:2014raa ; Behtash:2018voa ; Fujimori:2019skd ; Misumi:2019upg ; Fujimori:2020zka ; Unsal:2020yeh ; Pazarbasi:2021ifb ; Pazarbasi:2021fey . It also has deep connections with 4d ${\cal N}=2$ gauge theories Nekrasov:2009rc ; Mironov:2009uv ; Kashani-Poor:2015pca ; Kashani-Poor:2016edc ; Ashok:2016yxz ; Yan:2020kkb , wall-crossing phenomena Gaiotto:2012rg ; Allegretti:2020dyt , ODE/IM correspondence Dorey:2001uw ; Dorey:2007zx ; Ito:2018eon ; Ito:2019jio ; Imaizumi:2020fxf , TBA equations Emery:2020qqu ; Ito:2024nlt topological string theory Grassi:2014cla ; Grassi:2014zfa ; Codesido:2017dns ; Codesido:2017jwp ; Hollands:2019wbr ; Ashok:2019gee ; Coman:2020qgf ; Iwaki:2023cek and other subjects Kuwagaki:2020pry ; Taya:2020dco ; Enomoto:2020xlf ; Enomoto:2022mti ; Duan:2018dvj ; Imaizumi:2022dgj ; vanSpaendonck:2022kit ; Kamata:2023opn ; Kamata:2024tyb ; Bucciotti:2023trp ; vanSpaendonck:2024rin .
@@
An important property of the EWKB method is that it always fully encods the physical quantities with the associated resurgence structures. Therefore, by construction our approach links the resurgence structures of different sectors. To understand this relationship better, we show that the median summations of exact quantization conditions have the same form in all sectors with an appropriate redefinition of $A$-cycle, in a way that compatible with the analytic continuation of the associated quantum action. This manifestly displays the preservation of resurgence structures in all sectors as well as the smooth transition of the spectrum across the barrier top.
To further develop the EWKB method, we also extend the Weber-type approach. While the Airy-type approach to EWKB has been extensively studied for its qualitative insights, the Weber-type approach offers a more quantitative framework, particularly when dealing with quantum actions around saddle points. In this work, we generalize the Weber-type EWKB method to handle all types of saddle points, i.e. both minima and maxima of a given classical potential. In this way, we derive exact estimates for quantum actions, labeled as $a(u)$ and $a^{\mathrm{D}}(u)$, highlighting the local and non-local components of these actions.
Our analysis, also, reveals an important symmetry, i.e. S-duality, that governs the exchange between perturbative and non-perturbative behaviors around minima and maxima. S-duality is a central concept in many areas of theoretical physics, as it relates weak and strong coupling regimes in quantum systems. In Codesido:2017dns ; Codesido:2017jwp , this exchange between minima and maxima was discussed in the context of the relationship between holomorphic anomaly equations and all-order WKB expansions of one dimensional quantum systems. Here, we show that it is natural by product of the EWKB analysis and become manifest when the Weber-type approach is utilized. In addition to that the nature of EWKB analysis indicates that the S-duality, which is originally defined at the leading order (classical level), could relate the resurgence structures of dual theories. In this sense, we also conjecture a generic S-transformation of P-NP relations between dual theories.
@@
Finally, we analyze two Chebyshev potentials, periodic and symmetric double-well potentials in details. They are prototypical systems for studying tunneling phenomenon as well as multi-instanton contributions and resurgence theory Sueishi:2020rug ; Sueishi:2021xti ; Alvarez3 ; Zinn-Justin:2004qzw ; Zinn-Justin:2004vcw ; Dunne:2013ada ; Dunne:2014bca ; Misumi:2015dua ; Basar:2015xna . For our purposes, they serve as rich examples for understanding the evolution of quantum actions across different sectors while preserving the entire resurgence structure in the presence of multiple degenerate saddle points.
@@ 2 Generalization of the EWKB method
Note that despite the apparent differences of $P(x,u,g)$, the both setups are in fact equivalent to each other when the quantization of (1) is considered. The only difference in two cases is simply rescaling the parameter $u$ as $u\rightarrow u_{0}+g\tilde{u}$, which makes $\tilde{u}$ a dimensionless parameter, and prepares the setup for Weber-type analysis. We clarify the last point in Section 2.2.
This simple relationship between the Airy-type and Weber-type setups enables us to use either one in quantization of (1) and they provide equivalent but complementary information about the quantum system. On the one side, the Airy-type approach helps us to construct exact quantization conditions in terms of the quantum actions of the curve $P(x,u,g)$, which was discussed in detail in Sueishi:2020rug ; Sueishi:2021xti ; Kamata:2021jrs . For our purposes, the Airy-type approach is also crucial to explicitly understand the continuous transition between different sectors of a given classical potential $V(x)$. The Weber-type EWKB, on the other hand, is useful in a more quantitative manner as it provides explicit expressions for the quantum actions. For our purposes, the Weber-type approach helps us to identify the changes in the quantum actions upon a transition between different sectors. Moreover, in computations of the perturbative actions, the Weber-type setup provides a powerful tool to obtain the series expansions via a simple residue analysis.
We finally note that even though the Weber-type EWKB’s quantitative power, in computations of the non-perturbative part, we realize that Airy-type integrals becomes more convenient. So, when we discuss the details of the non-perturbative actions, which we obtain via the Weber-type approach, we return to the Airy-type integrals. This is allowed by virtue of the relationship between the two approaches which is carried to the WKB expansions Sueishi:2021xti .
@@ 2.1 Airy-type approach
The Airy-type approach is very well studied for the below-barrier-top sectors of $V(x)$ DDP2 ; Sueishi:2020rug ; Sueishi:2021xti ; Kamata:2021jrs and it was shown that it is in complete agreement with the other methods. In this subsection, we present a novel approach to uncover the transition to the above-barrier-top sectors. In order to be precise in our discussions, we only focus on the inverted harmonic oscillator and postpone the exact quantization of more generic potential to Sections 4 and 5. Note that despite the simplicity of the inverted harmonic oscillator, as we are interested in potentials with locally harmonic saddles, it provides a great approximation for the regions around these saddles. Therefore, our discussion in this section generalizes for generic locally harmonic potentials in a straightforward manner.
Since the setup of Airy-type EWKB is well-known and thoroughly discussed in the literature, we do not present it in details. However, in order to understand the transition between different sectors, we need to modify the setup slightly by introducing a complexification to the spectral parameter $u$. This complexification is analogous to $u\rightarrow u\pm i\varepsilon$ which we discuss in Section 3 and helps us to prevent from the singular structure at the top of a barrier [^1] Another singular point in Airy-type approach corresponds to the bottom of a well. Once $u$ reaches a singular point, the geometry of Stokes diagrams changes discontinously. The new type of Stokes diagrams subject of the Weber-type EWKB which we discuss in Section 2.2. while transitioning from one sector to another one.
@@
Inserting the ansatz in (6), we get the Ricatti equation
@@
$\tilde{s}=\sum_{m=0}^{\infty}s_{2m}\,g^{2m}\,,\qquad\hat{s}=\sum_{m=0}^{\infty}s_{2m-1}\,g^{2m-1}\,,$ (11)
@@
Note that in (14), we set the lower boundary of the integral to a (simple) turning point of the classical potential $V(x)$, i.e $P_{\mathrm{A}}(x_{0},u)=0$. This is known as a normalization point and it plays a crucial role in Airy-type EWKB analysis.
@@
Note that last matrix is due to the square root branch cut of the term $\sqrt{P_{\mathrm{A}}}$ and it implies exchange of the solutions $\psi^{\pm}$. While constructing the Stokes diagram of a given curve $P_{\mathrm{A}}$, the branch cuts are put by a chosen convention. See Fig. 1(b) for an example. We discuss on the equivalence of different branch cut choices in Appendix A.
@@
Finally, we introduce the matrix (also called Voros matrix) for changing the normalization point when there are more than one turning point
@@
When a potential $V(x)$ has two or more turning points, it is possible that a Stokes line emerging from a turning point ends at another one, which indicates the degeneracy of Stokes lines [^2] Note that a degenerate Stokes diagram is sometimes called as merging turning points in the literature it is also called merging turning points AKT4 ; Takei3 ; Sueishi:2021xti . In this paper, we stick with the term “degenerate” as merging turning points also means their coalescence at the saddle points of a given potential.. The degeneracy signals a need of analytic continuation [^3] More precisely, it means there is another non-Borel summable series. This is the perturbation series of $u$ and it should be promoted to trans-series by including non-trivial non-perturbative contributions. While the degeneration of a Stokes diagram means one of the series solutions for $\psi^{\pm}$ is non-Borel summable, it doesn’t mean that the associated perturbative series for $u$ is also non-Borel summable. One such example is the stable quartic oscillator. We refer to Kamata:2023opn for a detailed discussion. which deforms the Stokes diagram so that two turning points would not be connected by a Stokes line. In the literature, in general, the analytic continuation of the Stokes diagrams are handled by a deformation of the parameter $g$. For example, if the condition (16) is satisfied when $\arg g=0$, then introducing a small imaginary part to $g$ would deform the Stokes diagrams and break the degeneracy.
Figure 2: The inverted harmonic oscillator and associated turning points for $u<1$ (equivalently $\theta=0$) and $u>1$ (equivalently $\theta=\pm\pi$). Note that for $\theta=\pm\pi$, the locations of the turning points $x_{\pm}$ reversed, which indicates two different analytic continuations. (See Fig. 5 for its effect on the Stokes geometry.) This is also the case for generic locally harmonic potentials and we encounter with it when we discuss periodic and symmetric double-well potentials in Sections 4 and 5.
For our purposes, on the other hand, we analytically continue the parameter $u$ by introducing a phase as $u=|u|e^{i\alpha}$. This is equivalent to introducing an imaginary part to $g$. In order to see this, let us investigate $\sigma_{\mathrm{A}}$ for an inverted harmonic oscillator. Without losing any generality we set the origin at $x=0$ and $u=0$. Then, introducing a phase for the parameter $u$, we write $P_{\mathrm{A}}$ as
@@
where we introduced a phase for $g$. Equation (22) manifestly shows that phases $\theta_{u}$ and $\theta_{g}$ deforms the Stokes diagrams in opposite directions and both could be used to break the degeneracy of Stokes diagrams. For example around $\theta_{u}=0$, the deformation and degeneracy breaking of the Stokes diagram for the curve (21) is depicted in Fig. 3. Note that the geometries of the diagrams for $\theta_{u}=0^{\pm}$ are the same as the ones for $\theta_{g}=0^{\mp}$.
@@
Figure 4: The transition of Stokes lines associated with the turning points by the phase rotation $\theta_{u}=0\rightarrow\pm\pi$
Finally, we present the transition of the Stokes diagram of Fig. 2 from the sector $u<0$ to $u>0$. Two possible analytic continuations corresponding to the transitions $\theta_{u}=0^{+}\rightarrow\pi^{-}$ and $\theta_{u}=0^{-}\rightarrow-\pi^{+}$ are depicted in Fig. 5, which clearly shows the continuous transition between the two sectors since no Stokes line or branch cut become degenerate throughout the rotation. Note that although the Stokes lines degenerate at $\theta_{u}=\pm\pi$ in the same way, the difference in the branch cuts’ orientation indicates that two Stokes diagrams are not identical [^4] This becomes an important aspect of the analytic continuation we introduce in this section when we discuss more general potentials in Sections 4 and 5. We also elaborate the relationship between different branch cut conventions in Appendix A..
Figure 5: The analytic continuations of the Stokes diagram while the phase changing as $\theta_{u}=0\rightarrow\pm\pi$. Note that all the diagrams are originated from the same degenerate diagram at $\theta=0$ and two possible analytic continuations lead to different branch cut conventions at $\theta_{u}=\pm\pi$. Throughout the transitions, except $\theta=0$ and $\theta=\pm\pi$, there is no Stokes lines degeneration which implies the deformation is a continuous one.
@@
Having a single transition matrix for all sectors is also the case for all potentials with locally harmonic maxima. It leads to a single quantization condition for all sectors for any potentials we discuss in Sections 4-5. Before the discussion of individual potentials, in the next subsection, we first present the Weber type EWKB which provides us a quantitative picture.

### v2

@@ Abstract
We explore the exact-WKB (EWKB) method through the analysis of Airy and Weber types, with an emphasis on the exact quantization of locally harmonic potentials in multiple sectors. The core innovation of our work lies in introducing a novel complexification approach to the energy parameter $u$, distinct from the common complexification of the (semi-classical) expansion parameter used in Borel summability. This new technique allows for continuous analytical continuation across different sectors of a potential while maintaining the exact quantization condition, even before median summation. By redefining the $A$-cycle above the potential barrier top, we ensure the quantization condition remains real and, by use of the Stokes automorphism and the median resummation, show that the resurgence structure is preserved across transitions between sectors. Furthermore, we discuss the Weber-type exact-WKB method, offering exact estimates for quantum actions around all types of saddle points, generalizing previous results. Through the analysis of these quantum actions, we reveal the presence of $S$-duality, facilitating the exchange between perturbative and non-perturbative behaviors, and we conjecture the mapping of the P-NP relations between dual theories. Our study encompasses periodic and symmetric double-well potentials, demonstrating that the exact-WKB method captures intricate structures in quantum systems in all sectors, including multi-instanton contributions and the resurgence of quantum actions.
@@ 1 Introduction
The exact-WKB (Wentzel-Kramers-Brillouin) method BPV ; Voros1983 ; Silverstone ; AKT1 ; AKT2 ; Takei1 ; DDP2 ; DP1 ; Takei2 ; Kawai1 ; Takei3 ; AKT3 ; AKT4 ; Schafke1 ; Iwaki1 stands as one of the most powerful tools in the study of quantum mechanical systems. With being combined with Borel resummation, the exact-WKB (EWKB) method offers significant improvements over semi-classical approximations, particularly in quantum systems where both perturbative and non-perturbative effects play a role. In recent decades, the EWKB method has seen significant advancements, fueled by contributions from both physicists and mathematicians. These developments have led to a deeper understanding of the connection between the semi-classical analysis and the exact quantization conditions Sueishi:2020rug ; Sueishi:2021xti ; Kamata:2021jrs , revealing intricate structures such as Stokes phenomena and resurgence theory in quantum mechanical systems Alvarez1 ; Alvarez2 ; Alvarez3 ; Alvarez4 ; Jentschura:2004jg ; Zinn-Justin:2004vcw ; Zinn-Justin:2004qzw ; Jentschura:2010zza ; Jentschura:2011zza ; Basar:2013eka ; Dunne:2013ada ; Dunne:2014bca ; Basar:2015xna ; Misumi:2015dua ; Behtash:2015loa ; Gahramanov:2015yxk ; Kozcaz:2016wvy ; Dunne:2016jsr ; Dunne:2016qix ; Fujimori:2016ljw ; Basar:2017hpr ; Fujimori:2017oab ; Sueishi:2019xcj ; Cavusoglu:2023bai ; Cavusoglu:2024usn ; Ture:2024nbi . The method has been instrumental in exploring non-perturbative effects involving multi-instanton contributions which appear in analysis of various quantum mechanical and field theoretical systems Zinn-Justin:1981qzi ; Zinn-Justin:1983yiy ; Unsal:2007vu ; Unsal:2007jx ; Shifman:2008ja ; Poppitz:2009uq ; Anber:2011de ; Poppitz:2012sw ; Misumi:2014raa ; Behtash:2018voa ; Fujimori:2019skd ; Misumi:2019upg ; Fujimori:2020zka ; Unsal:2020yeh ; Pazarbasi:2021ifb ; Pazarbasi:2021fey . It also has deep connections with 4d ${\cal N}=2$ gauge theories Nekrasov:2009rc ; Mironov:2009uv ; Kashani-Poor:2015pca ; Kashani-Poor:2016edc ; Ashok:2016yxz ; Yan:2020kkb , wall-crossing phenomena Gaiotto:2012rg ; Allegretti:2020dyt , ODE/IM correspondence Dorey:2001uw ; Dorey:2007zx ; Ito:2018eon ; Ito:2019jio ; Imaizumi:2020fxf , TBA equations Emery:2020qqu ; Ito:2024nlt topological string theory Grassi:2014cla ; Grassi:2014zfa ; Codesido:2017dns ; Codesido:2017jwp ; Hollands:2019wbr ; Ashok:2019gee ; Coman:2020qgf ; Iwaki:2023cek and other subjects Kuwagaki:2020pry ; Taya:2020dco ; Enomoto:2020xlf ; Enomoto:2022mti ; Duan:2018dvj ; Imaizumi:2022dgj ; vanSpaendonck:2022kit ; Kamata:2023opn ; Kamata:2024tyb ; Bucciotti:2023trp ; vanSpaendonck:2024rin .
@@
An important property of the EWKB method is that it always fully encodes the physical quantities with the associated resurgence structures. Therefore, by construction our approach links the resurgence structures of different sectors. To understand this relationship better, we show that the median summations of exact quantization conditions have the same form in all sectors with an appropriate redefinition of $A$-cycle, in a way that compatible with the analytic continuation of the associated quantum action. This manifestly displays the preservation of resurgence structures in all sectors as well as the smooth transition of the spectrum across the barrier top.
To further develop the EWKB method, we also discuss the Weber-type approach. While the Airy-type approach to EWKB has been extensively studied for its qualitative insights, the Weber-type approach offers a more quantitative framework, particularly when dealing with quantum actions around saddle points. In this work, we generalize the Weber-type EWKB method to handle all types of saddle points, i.e. both minima and maxima of a given classical potential. Such a generalization was previously discussed in DDP2 ; DP1 from a mathematical point of view where authors provided double-well potential as an explicit example. In our discussion, we follow a different path, which is based on AKT1 ; AKT4 , in the construction of the Weber-type EWKB. Moreover, we provide exact expressions of quantum actions, labeled as $a(u,g)$ and $a^{\mathrm{D}}(u,g)$, for generic potentials, highlighting the local and non-local components of these actions as well as their meaning in the physical literature.
In addition to that providing exact estimates for generic potentials, our analysis reveals an important symmetry, i.e. $S$-duality, that governs the exchange between perturbative and non-perturbative behaviors around minima and maxima. $S$-duality is a central concept in many areas of theoretical physics, as it relates weak and strong coupling regimes in quantum systems. In Codesido:2017dns ; Codesido:2017jwp , this exchange between minima and maxima was discussed in the context of the relationship between holomorphic anomaly equations and all-order WKB expansions of one dimensional quantum systems. Here, we show that it is a natural by product of the EWKB analysis and become manifest when the Weber-type approach is utilized. In addition to that the nature of EWKB analysis indicates that the $S$-duality, which is originally defined at the leading order (classical level), could relate the resurgence structures of dual theories. In this sense, we also conjecture a generic $S$-transformation of P-NP relations between dual theories.
@@
Finally, we discuss two Chebyshev potentials, periodic and symmetric double-well potentials in details. They are prototypical systems for studying tunneling phenomenon as well as multi-instanton contributions and resurgence theory Sueishi:2020rug ; Sueishi:2021xti ; Alvarez3 ; Zinn-Justin:2004qzw ; Zinn-Justin:2004vcw ; Dunne:2013ada ; Dunne:2014bca ; Misumi:2015dua ; Basar:2015xna . For our purposes, they serve as rich examples for understanding the evolution of quantum actions across different sectors while preserving the entire resurgence structure in the presence of multiple degenerate saddle points.
@@ 2 Generalization of the EWKB method
Note that despite the apparent differences in $P(x,u,g)$, the both setups are in fact equivalent to each other when the quantization of (1) is considered. The only difference in two cases is simple rescaling of the parameter $u$ as $u\rightarrow u_{0}+g\tilde{u}$, which makes $\tilde{u}$ a dimensionless parameter, and prepares the setup for Weber-type analysis. We clarify the last point in Section 2.2.
The relationship between the Airy-type and Weber-type setups enables us to use either one in the quantization of (1) and they provide equivalent but complementary information about the quantum system. On the one side, the Airy-type approach helps us to construct exact quantization conditions in terms of the quantum actions of the curve $P(x,u,g)$, which was discussed in detail in Sueishi:2020rug ; Sueishi:2021xti ; Kamata:2021jrs . For our purposes, the Airy-type approach is also crucial to explicitly understand the continuous transition between different sectors of a given classical potential $V(x)$. The Weber-type EWKB, on the other hand, is useful in a more quantitative manner as it provides explicit expressions for the quantum actions. For our purposes, the Weber-type approach helps us to obtain the explicit expressions for the quantum actions in all sectors and establish the relationship between the actions in different sectors which indicates an $S$-duality from the EWKB point of view.
We finally note that despite the Weber-type EWKB’s quantitative power, in computations of the non-perturbative actions, we realize that Airy-type integrals becomes more convenient. So, when we discuss the details of the non-perturbative actions, which we obtain via the Weber-type approach, we return to the Airy-type integrals. This is allowed by virtue of the relationship between the two approaches which is carried to the WKB expansions DP1 ; Sueishi:2021xti .
@@ 2.1 Airy-type approach
The Airy-type approach is very well studied for the below-barrier-top sectors of $V(x)$ DDP2 ; Sueishi:2020rug ; Sueishi:2021xti ; Kamata:2021jrs and it was shown that it is in complete agreement with the other methods. In this subsection, we present a novel approach to uncover the transition to the above-barrier-top sectors. In order to be precise in our discussions, we only focus on the inverted harmonic oscillator and postpone the exact quantization of more generic potential to Sections 4 and 5. Note that despite the simplicity of the inverted harmonic oscillator, since we are interested in potentials with locally harmonic saddles, it provides a great approximation for the regions around these saddles. Therefore, our discussion in this section generalizes for generic locally harmonic potentials in a straightforward manner.
Since the setup of Airy-type EWKB is well-known and thoroughly discussed in the literature, we do not present it in details. However, in order to understand the transition between different sectors, we need to modify the setup slightly by introducing a complexification to the spectral parameter $u$. Note that this complexification is analogous to $u\rightarrow u\pm i\varepsilon$ which we discuss in Section 3 and helps us to prevent from the singular structure at the top of a barrier [^1] Another singular point in Airy-type approach corresponds to the bottom of a well. Once $u$ reaches a singular point, the geometry of Stokes diagrams changes discontinously. The new type of Stokes diagrams subject of the Weber-type EWKB which we discuss in Section 2.2. while transitioning from one sector to another one.
@@
Inserting the ansatz in (6), we get the Riccati equation
@@
$\hat{s}=\sum_{m=0}^{\infty}s_{2m}\,g^{2m}\,,\qquad\tilde{s}=\sum_{m=0}^{\infty}s_{2m-1}\,g^{2m-1}\,,$ (11)
@@
Note that in (14), we set the lower boundary of the integral to a (simple) turning point of the classical potential $V(x)$, i.e $P_{\mathrm{A}}(x_{0},u)=0$, which is also known as a normalization point.
@@
Note that last matrix is due to the square root branch cut of the term $\sqrt{P_{\mathrm{A}}}$ and it governs the exchange of the solutions $\psi^{\pm}$. While constructing the Stokes diagram of a given curve $P_{\mathrm{A}}$, the branch cuts are put by a chosen convention. See Fig. 1(b) for an example. We discuss the equivalence of different branch cut choices in Appendix A.
@@
Finally, we introduce a matrix (also called Voros matrix) for changing the normalization point when there are more than one turning point
@@
When a potential $V(x)$ has two or more turning points, it is possible that a Stokes line emerging from a turning point ends at another one, which indicates the degeneracy of Stokes lines [^2] Note that a degenerate Stokes diagram is sometimes called as merging turning points in the literature AKT4 ; Takei3 ; Sueishi:2021xti . In this paper, we stick with the term “degenerate” as merging turning points also means their coalescence at the saddle points of a given potential, which is the subject of the Weber-type EWKB approach.. The degeneracy signals a need of analytic continuation [^3] Different analytic continuations which breaks the degeneracy related to each other via so-called DDP formula DDP2 ; DP1 which is deeply related to Borel summation process Kamata:2023opn . We utilize it in Sections 4 and 5 to obtain the median exact quantization conditions which reveals manifestly real spectra. which deforms the Stokes diagram so that two turning points would not be connected by a Stokes line. In the literature, in general, the analytic continuation of the Stokes diagrams are handled by a deformation of the parameter $g$. For example, if the condition (16) is satisfied when $\arg g=0$, then introducing a small imaginary part to $g$ would deform the Stokes diagrams and break the degeneracy.
Figure 2: Inverted harmonic oscillator and associated turning points for $u<0$ (equivalently $\theta_{u}=0$) and $u>0$ (equivalently $\theta_{u}=\pm\pi$). Note that for $\theta_{u}=\pm\pi$, the locations of the turning points $x_{\pm}$ reversed, which indicates two different analytic continuations. (See Fig. 5 for its effect on the Stokes geometry.) This is also the case for generic locally harmonic potentials and we encounter with it when we discuss periodic and symmetric double-well potentials in Sections 4 and 5.
For our purposes, on the other hand, we analytically continue the parameter $u$ by introducing a phase as $u=|u|e^{i\theta_{u}}$. This is equivalent to introducing an imaginary part to $g$. In order to see this, let us investigate $\sigma_{\mathrm{A}}$ for an inverted harmonic oscillator. Without losing any generality we set the origin at $x=0$ and $u=0$. Then, introducing a phase for the parameter $u$, we write $P_{\mathrm{A}}$ as
@@
where we introduced a phase for $g$. Equation (22) manifestly shows that phases $\theta_{u}$ and $\theta_{g}$ deform the Stokes diagrams in opposite directions and both could be used to break the degeneracy of Stokes diagrams. For example around $\theta_{u}=0$, the deformation and degeneracy breaking of the Stokes diagram for the curve (21) is depicted in Fig. 3. Note that the geometries of the diagrams for $\theta_{u}=0^{\pm}$ are the same as the ones for $\theta_{g}=0^{\mp}$.
@@
Figure 4: The transition of Stokes lines associated with the turning points $x=x_{\pm}$ by the phase rotation $\theta_{u}=0\rightarrow\pm\pi$.
Finally, we present the transition of the Stokes diagram of Fig. 2 from the sector $u<0$ to $u>0$. Two possible analytic continuations corresponding to the transitions $\theta_{u}=0^{+}\rightarrow\pi^{-}$ and $\theta_{u}=0^{-}\rightarrow-\pi^{+}$ are depicted in Fig. 5, which clearly shows the continuous transition between the two sectors since no Stokes line or branch cut become degenerate throughout the rotation. Note that although the Stokes lines are again degenerated at $\theta_{u}=\pm\pi$ in the same way, the difference in the branch cuts’ orientation indicates that two Stokes diagrams are not identical [^4] This becomes an important aspect of the analytic continuation we introduce in this section, when we discuss more general potentials in Sections 4 and 5. We also elaborate the relationship between different branch cut conventions in Appendix A..
Figure 5: The analytic continuations of the Stokes diagram while the phase changing as $\theta_{u}=0\rightarrow\pm\pi$. Note that all the diagrams are originated from the same degenerate diagram at $\theta_{u}=0$ and two possible analytic continuations lead to different branch cut conventions at $\theta_{u}=\pm\pi$. Throughout the transitions, except $\theta_{u}=0$ and $\theta_{u}=\pm\pi$, there is no Stokes lines degeneration which implies the deformation is a continuous one.
@@
Having a single transition matrix for all sectors is also the case for all potentials with locally harmonic maxima and it leads to a single quantization condition for all sectors for any potentials with degenerate saddle points Before discussing the details of the transition, in the next subsection, we first present the Weber type EWKB which provides us a quantitative picture.